Over the past 5 years, NBA's common stock earnings per share have grown from $0.62 to...

  1. Over the past 5 years, NBA's common stock earnings per share have grown from $0.62 to $0.91. If an investor is NBA stock is assumed to have a required rate of return of 14%, what is the estimated value of NBA if its current dividend is $0.12? Assume dividends will continue to grow at a rate similar to that of EPS.

    A.

    $2.16      

    B.

    $1.62

    C.

    $4.94

    D.

    $2.00
Answer #1

FV = PV (1+g)^n

0.91 = 0.62 ( 1 + g)^5

(1+g)^5 = 0.91 / 0.62

= 1.4677

1+g = 1.4677^(1/5)

= 1.0798

g =1.0798 - 1

= 0.0798 i.e 7.98%

P0 = D0(1+g) / [ Ke - g ]

= $ 0.12 ( 1 + 0.0798 ) / [ 14% - 7.98% ]

= 0.12 * 1.0798 / [ 6.02% ]

= 0.1296 / 6.02%

= 2.16

OPtion A is correct.